### v1

# CNSocialDepress: A Chinese Social Media Dataset for Depression Risk Detection and Structured Analysis

Abstract
Depression is a pressing global public health issue, yet publicly available Chinese-language resources for risk detection remain scarce and are mostly limited to binary classification. To address this limitation, we release CNSocialDepress, a benchmark dataset for depression risk detection from Chinese social media posts. The dataset contains 44,178 texts from 233 users, within which psychological experts annotated 10,306 depression-related segments. CNSocialDepress provides binary risk labels together with structured multi-dimensional psychological attributes, enabling interpretable and fine-grained analysis of depressive signals. Experimental results demonstrate its utility across a wide range of NLP tasks, including structured psychological profiling and fine-tuning of large language models for depression detection. Comprehensive evaluations highlight the dataset’s effectiveness and practical value for depression risk identification and psychological analysis, thereby providing insights to mental health applications tailored for Chinese-speaking populations.
CNSocialDepress: A Chinese Social Media Dataset for Depression Risk Detection and Structured Analysis
Jinyuan XU1,*, Tian LAN2,*, Xintao YU3, Xue HE3,4, Hezhi ZHANG5, Ying WANG6, Pierre Magistry1, Mathieu Valette1, Lei LI7,8†thanks: * These authors contributed equally to this work.thanks: † Corresponding Author. 1 Ertim Inalco, 2 Milkuya Studio, 3 Sorbonne Université, 4 IRD Lab 5 Faculty of Psychology, Peking University 6 Faculty of Psychology and Cognitive Science, Beijing Normal University 7 University of Washington, 8 VitaSight
1 Introduction
Depressive disorders are among the most common mental health conditions worldwide, characterized by persistent low mood or loss of interest in daily activities. According to a 2023 report [^1] https://www.who.int/news-room/fact-sheets/detail/depression by the World Health Organization (WHO), approximately 280 million people globally suffer from depression. In China, a 2024 survey by the Chinese Center for Disease Control and Prevention Wang et al. (2024a) estimates that around 95 million individuals are affected by depression, and out of the 280,000 suicides reported annually, 40% are linked to depressive disorders. Moreover, research Wang et al. (2024a) shows that depression strongly correlates with both suicidal and non-suicidal self-injurious deaths.
Motivated by the urgent need for the detection of depression, researchers increasingly apply machine learning (ML) and natural language processing (NLP) methods for automated analysis of depression risk Squires et al. (2023); Hasib et al. (2023); Aleem et al. (2022). While early efforts have shown promising results, they are heavily constrained by the limitations of existing datasets. Traditional depression detection studies often rely on clinical dataBittar et al. (2019); Fernandes et al. (2018), or on transcripts from medical or psychological interviews resources Shen et al. (2022); Li et al. (2022) that are are expensive to collect, limited in size and diversity, and may not reflect the informal, emotionally nuanced expressions typical of real-world online environments.
To overcome these limitations, recent research has shifted toward leveraging user-generated content on social media platforms [^2] https://github.com/bucuram/depression-datasets-nlp Harrigian et al. (2021), which provides rich linguistic signals and greater availability. Such data are also more easily anonymized, alleviating some privacy concerns. Beyond these methodological advantages, social media itself has become an emotional outlet for many individuals, especially in East Asian contexts Yang and Li (2009); Zhou et al. (2023), where personal emotions are more freely expressed online rather than in face to face settings. Social media serves as a critical outlet for emotional expression, potentially offering richer and more authentic linguistic data for depression analysis.
However, most existing datasets are in English and focus solely on classification with binary or multi-class labels, lacking structured psychological insight or professional validation. Cai et al. (2023); Wang et al. (2020); Li et al. (2020). Moreover, with the emergence of generative models applied to mental health Hu et al. (2024); Xu et al. (2024); Lai et al. (2023); Yang et al. (2024), many scholars in psychology and computational linguistics now emphasize the need for datasets combining risk labels with structured analyses, such as user profiles or survey-based explanations. Such datasets align more closely with real-world settings, where multi-faceted insights are helpful for clinical intervention.
Some recent efforts have introduced summarization style datasets for depression, leveraging information extraction and automated summarization techniques Sotudeh et al. (2022, 2021). However, these datasets usually rely on model generated content validated by automatic metrics (e.g., ROUGE Lin (2004)) and typically lack professional mental health annotation, raising concerns about domain-specific reliability and accuracy.
To address these challenges, we introduce (CNSocialDepress, the first publicly available Chinese-language dataset for depression risk detection that integrates binary risk labels with structured psychological analyses. All annotations are drafted and validated by certified mental health experts, ensuring both domain relevance and annotation quality. CNSD is designed to support multiple task paradigms, including binary classification, generation of structured analyses, summarization, and fine-tuning of large language models (LLMs) for psychological reasoning.
Our contributions are fourfold:
  • We construct a high-quality Chinese dataset that bridges the gap between classification and interpretable analysis in depression detection, and supports a wide range of tasks including binary classification, generation of structured psychological analyses, summarization, and fine-tuning of large language models.
  • We design a robust annotation and generation pipeline involving professional experts and structured templates;
  • We conduct extensive experiments demonstrating that CNSD enables effective training and evaluation of LLM-based methods across diverse depression-related tasks.
  • We propose and validate a robust pipeline for generating structured psychological analyses in the context of depression risk (Section 4).
2 Related Work
2.1 Existing Datasets for Depression Detection
Current depression detection datasets primarily come from English-language social media platforms like Twitter, Reddit, Facebook, and Instagram Shen et al. (2017); Parapar et al. (2022); Zhang et al. (2021); Raihan et al. (2024). Annotation methods include self-disclosure identification Yates et al. (2017); Bathina et al. (2021); Islam et al. (2018), manual coding by clinical raters Almouzini et al. (2019); Yazdavar et al. (2020); Alhamed et al. (2024), and symptom mapping Seabrook et al. (2018); Aldarwish and Ahmad (2017); Zhang et al. (2022) based on DSM American Psychiatric Association et al. (2000) and PHQ Kroenke et al. (2001). While most datasets use binary classification, others incorporate severity scales or symptom-specific tagging Zhang et al. (2022); Mowery et al. (2017). Multilingual resources for depression detection exist for Spanish Romero et al. (2024), Arabic Maghraby and Ali (2022), Russian Stankevich et al. (2020), Portuguese Santos et al. (2024), Japanese Yuka Niimi (2021), and Thai Hämäläinen et al. (2021). For Chinese, Sina Weibo is the primary source Li et al. (2020); Shen et al. (2018); Li et al. (2023); Guo et al. (2023); Yang et al. (2021), with datasets like WU3D Wang et al. (2020) and SWDD Cai et al. (2023) commonly used.
2.2 Methodological Advancements
Early methods in depression detection used statistical approaches like TF-IDF and feature engineering Yang et al. (2020) along with traditional machine learning classifiers Cortes (1995); Li (2024); Breiman (2001); McCallum et al. (1998); Dreiseitl and Ohno-Machado (2002); Li et al. (2024); Shi et al. (2024). With the rise of neural networks Schmidhuber (2015), deep learning models such as Word2Vec Mikolov et al. (2013), LSTM Sak et al. (2014), CNN Kim (2014), Transformer Vaswani (2017), and BERT Devlin et al. (2019) improved the detection of depression. Recently, large language models (LLMs) have been applied to depression detection tasks due to their ability to generate labels and provide interpretable analyses Lan et al. (2024); Hu et al. (2024); Xin Yan (2023); Lai et al. (2023). These models enhance explainability in mental health analysis Wang et al. (2024c); Yang et al. (2024); Xu et al. (2024); Yang et al. (2023); Hu et al. (2024); Xu et al. (2024).
2.3 Summary and Gaps
Despite existing efforts, current datasets rarely include structured psychological analysis, especially for non-English social media content. Furthermore, available resources often lack professional mental-health annotation, limiting their practical utility. This study addresses these limitations by presenting the first expert-validated Chinese-language dataset featuring detailed structured annotations, thereby providing richer resources for depression-risk detection and analysis.
3 Construction of Dataset
Figure 1: Dataset Construction Process: During the data annotation process, we used a subset of the original SWDD dataset for labeling, which included 116 Positive (Depressive) users and 117 Negative users. Psychologists used rating scales based on the DSM-5 and the statistical results of the dataset’s text to formulate an initial labeling guideline. Then, during the psychologists’ labeling process, random sampling was continuously performed for cross validation, and the labeling standards were continuously updated based on the annotation outcomes. Ultimately, the labeled gold-standard data was obtained, with each user containing a six-dimensional structural analyses summary.
Our raw data are derived from the SWDD dataset Cai et al. (2023), a user-level dataset collected from Sina Weibo, the world’s largest Chinese-language social media platform. This dataset comprises two categories of users (depressed and non-depressed), each user has between dozens and hundreds of Weibo posts. The dataset, which is labeled as either depressed or non-depressed, also contains user-level depression feature annotations provided by experts in accordance with the DSM-5333https://www.psychiatry.org/psychiatrists/practice/dsm/educational-resources/dsm-5-fact-sheets American Psychiatric Association et al. (2000) reference standards.
From this SWDD dataset, we have 116 self-reported depressed (positive) users and 117 non-depressed (negative) users to form the candidate samples for our dataset. Each selected user has at least 60 posts, and the total token count for each user’s posts is no fewer than 3,000 tokens [^4] Qwen2.5’s tokenizer.
To guide the annotation process, a group of psychology experts formulated a preliminary annotation standard based on the DSM-5 diagnostic criteria, the PHQ-9 scale Kroenke et al. (2001), and prior research on linguistic features indicative of depression in online texts Mothe et al. (2022).
Criteria involving objective factual statements (e.g., clinical symptoms, medical records, self-reports) with stronger diagnostic specificity are designated as Primary Criteria, while criteria based on subjective linguistic or emotional expressions (e.g., negative phrasing, emotional intensity) with weaker specificity are designated as lower-weight Secondary Criteria, and the hierarchical classification follows clinical diagnostic principles that prioritize objective symptoms over subjective perceptions. This standard comprises six depression dimensions:
  • Dimension 1: Depressive Psychological State (primary criterion). At its core, this category reflects a loss of self-worth, overwhelming guilt, and suicidal ideation. Representative semantic segments include Inferiority, Sorry, Death, Self-harm, Wrist cutting, Suicide, etc.
  • Dimension 2: Depression-Related Medical Expressions (primary criterion). Focusing on expressions linked to medication usage and psychological diagnoses, this category covers terms such as Take medication, Venlafaxine, Side effects, Depression, Anxiety disorder, Hospital, Doctor, Disease, etc.
  • Dimension 3: Depression-Related Clinical Symptoms (primary criterion). Addressing physiological and somatic alterations, including changes in weight or appetite, disturbed sleep patterns, fatigue, and various forms of physical pain, this dimension includes semantic segments such as Appetite, Insomnia, Sleeping pills, Nightmare, Tired, Headache, Stomachache, among others.
  • Dimension 4: Negative Emotions (secondary criterion). Encompassing a broad array of adverse emotional states, this part presents examples such as Torment, Abyss, Hell, Sad, Grieve, Anxious, Depressed, Lonely, Despair, etc.
  • Dimension 5: Potential External Causes of Depression (secondary criterion). Highlighting possible triggers like marital or intimate relationship issues, family conflicts, significant social events, and everyday stressors, this segment comprises semantic units including Divorce, Parents, College entrance examination, School, Teacher, Poor, Drop out, etc.
  • Dimension 6: Depression-Related Patterns of Language Use (secondary criterion). Emphasizing particular linguistic patterns, such as negations, interrogatives, and derogatory expressions. This category includes illustrative semantic segments like I don’t know, I don’t like me, I really don’t want to, I can’t do it, why, and what should I do?, among others.
Figure 2: Example entry from the CNSD-Gold dataset
The annotation work was carried out by four senior researchers, each with substantial research experience in psychology, specializing in depression scales. Using an internally developed annotation platform, they identified specific depression dimensions present in each post, along with the corresponding semantic segments. Throughout this process, the annotators periodically performed random cross-validation to ensure reliability and consistency. Because this annotation task relies on qualitative rather than quantitative metrics, the annotation schema was updated as needed, informed by real-time feedback from the annotated texts and discussions among experts. During the annotation process, the expert annotators did not refer to the binary labels from the original dataset, consequently, some of the depression risk labels may differ from those in the original data set.
This rigorous process produced the CNSD Gold dataset (100 positive and 100 negative users) and the CNSD Test dataset (16 positive and 17 negative users). The dataset construction process can be referenced in Figure 1.
Each entry as shown in Figure 2 comprises two parts: a binary depression label and a detailed, dimension-specific analysis. For each dimension, the analysis includes an overall assessment and a list of post numbers with brief justifications for their selection.
The overall statistics are presented in Table 1, with a total of 10,306 dimension-level annotations summarized in Table 2.
Positive (Depressive) Negative (Non-depressive)
Number of Users 116 117
Number of Texts 20,360 23,818
Number of Tokens 1,024,978 1,115,951
Table 1: Dataset Statistics.
Dimension Negative User Positive User
Dim1 117 2127
Dim2 2 555
Dim3 126 768
Dim4 514 2933
Dim5 193 863
Dim6 231 1877
Total 1183 9123
Table 2: Dimension Annotation Statistics for Negative and Positive Users: This table shows the statistics of depression repeated segments marked by psychological experts across different dimensions. A text may contain multiple depression repeated segments.
4 Automated Dataset Generation Pipeline
Due to the substantial human effort and professional psychological expertise required for dataset annotation, user-level depression risk datasets based on social media remain scarce in the Chinese domain. To address this gap, we developed an automated dataset generation pipeline. This pipeline was used to create the CNSD Silver dataset by processing the SWDD binary classification dataset, leveraging insights from the manually annotated CNSD Gold dataset (described in Section 3). For the construction of CNSD Silver, we randomly selected 100 positive (depressive risk) users and 100 negative (non-depressive risk) users from the SWDD dataset. Each user contributed between several dozen and several hundred posts, totaling more than 3,000 tokens per user.
The pipeline consists of two main modules:
4.1 Module I: Dimension-Wise Automatic Labeling
This module uses a smaller model (Qwen2.5-14B Bai et al. (2023)) to automatically annotate texts across six depression-related dimensions ($\mathcal{D}=\{D_{1},D_{2},\dots,D_{6}\}$). The model is first fine-tuned on carefully curated text-level data derived from the positive users in CNSD Gold, enabling it to assign each individual text to one or more of these six dimensions.
Data Preparation for Fine-tuning Module I Model:
  1. 1.
    Texts from positive users in CNSD Gold containing depression-related semantic segments for each dimension $D_{k}$ were extracted, forming sets $S_{k}$.
  2. 2.
    The dimension with the fewest samples ($n_{\text{min}}=442$, corresponding to "depression-related medical expressions") was identified:
    $n_{\text{min}}=\min_{k\in\{1,2,\dots,6\}}n_{k}=442$
  3. 3.
    To balance data across dimensions, texts for each dimension $D_{k}$ were downsampled to $n_{\text{min}}$ if $n_{k}>n_{\text{min}}$, forming balanced sets $S^{\prime}_{k}$:
    $S^{\prime}_{k}=\begin{cases}S_{k},&\text{if }n_{k}=n_{\text{min}},\\ \text{RandomSubset}(S_{k},n_{\text{min}}),&\text{if }n_{k}>n_{\text{min}}.\end{cases}$
    This resulted in $6\times 442$ labeled positive texts. Each text is annotated in the format: “This text segment belongs to [category] because it references [feature],” supplemented by 20 randomly generated expressions sharing the same meaning (see Appendix A.2).
  4. 4.
    For negative data, posts from negative users in CNSD Gold were split, and texts with any depression-related dimension were filtered out. From the remaining unlabeled texts, $6\times 442$ were uniformly sampled and assigned one of 20 distinct label expressions indicating absence of negative emotions (details in Appendix A.3).
  5. 5.
    The Module I model (Qwen2.5-14B) was then fine-tuned using LoRA Hu et al. (2022) on this combined set of $2\times(6\times 442)$ texts for 1 epoch at a learning rate of 5e-05 (details in Appendix A.4).
4.2 Module II: Automatic Verification and Summarization
In the second module, we applied the fine-tuned model from Module I to label every text for each user, extracting all texts that contain any depression-related semantics. We then leveraged the Deepseek R1 671B model DeepSeek-AI et al. (2025) to verify the correctness of both labels and texts on a case-by-case basis. Following this verification step, we employed a few-shot approach to summarize the extracted texts. In the few-shot setting, we instruct the model to assume the role of experts in psychology and text sentiment analysis. Additionally, we provide the model with detailed task requirements and exemplary gold-standard answer samples. Details are in the Appendix 3.
5 Experiments
We design our experiments around two central research questions. First , how effectively our data generation pipeline in producing high-quality structured annotations? Second , how well do various large language models (LLMs) perform on tasks involving user-level depression risk summarization and structured psychological analysis?
Based on these questions, our experiments aim to:
  1. 1.
    To assess the quality of text generated for six-dimensional depression analysis by models fine-tuned on CNSD Gold and CNSD Silver, comparing them against baseline models and few-shot approaches.
  2. 2.
    To evaluate the performance of these fine-tuned models on the downstream task of user-level depression risk classification.
  3. 3.
    To benchmark the performance of various leading LLMs on a structured depression analysis summarization task using our CNSD Gold dataset.
Additionally, as shown in the Appendix 5.5, we demonstrate the use of the CNSD Gold dataset as the test set for the binary depression classification task, further illustrating the dataset’s versatility.
5.1 Baseline Models
In the following experiments, we employ DeepSeek-R1 Distill-14B DeepSeek-AI et al. (2025), Qwen2.5-14B QwenTeam (2024), GPT-4o  [^5] https://openai.com/index/hello-GPT-4o/, GPT-4o-Mini [^6] https://openai.com/index/GPT-4o-mini-advancing-cost-efficient-intelligence/, DeepSeek-R1-671B [^7] https://www.deepseek.com/ DeepSeek-AI et al. (2025), and Llama3-8B-Chinese-Chat Wang et al. (2024b) as our generative baseline models. We summarize the key characteristics of each model on A.1.
5.2 Experimental Setup
For our experiments, we used NVIDIA A800 80G and A100 80G GPUs. Unless otherwise specified, all generation experiments were conducted with a temperature setting of 0.7. All fine-tuning experiments were conducted using the LLama Factory framework Zheng et al. (2024), employing the LoRA method Hu et al. (2022) for one epoch with a learning rate of 5e-5.
Our experiments consist of two main parts. The first part evaluates the performance of our pipeline on data generation, while the second part presents benchmarks on CNSD Gold dataset for Structured Summarization and Analysis Generation.
In our classification tasks, we use accuracy, recall, precision, and F1-score as evaluation metrics. For text generation experiments, we employ BLEU Papineni et al. (2002), ROUGE-1 Lin (2004), and BERTScoreZhang et al. (2020) for performance assessment.
5.3 Task I: Data Generation
In this section, we conducted multiple experiments on the CNSD Test dataset. We primarily evaluated the Qwen2.5 14B models fine-tuned on CNSD Gold and CNSD Silver:
  • Qwen2.5 14B Gold: Fine-tuned using the CNSD Gold.
  • Qwen2.5 14B Silver: Fine-tuned using the CNSD Silver. which was generated by applying the pipeline described in Section 4 to 100 randomly selected positive users and 100 randomly selected negative users from the SWDD binary classification dataset.
We compared the performance of these models with other baselines in two aspects:
  1. 1.
    The performance of the depression risk classification.
  2. 2.
    The quality of the generated text in the six-dimensional analyses section (including text hallucinations).
Throughout these experiments, unless noted otherwise, we employed a user-level generation strategy. This approach involves providing the complete set of a user’s Weibo posts along with an instruction that requires the model to directly output the user’s depression label along with the corresponding justification.
Strategy BERTScore ROUGE-1 BLEU
Pipeline 0.791 0.478 0.288
FS:Qwen2.5 14B 0.649 0.076 0.075
FS:GPT-4o 0.678 0.269 0.094
FS:GPT-4o Mini 0.674 0.170 0.070
FS:DeepSeek R1 671B 0.678 0.301 0.054
FS:DeepSeek R1 Distill -14B 0.6756 0.191 0.073
Table 3: Comparison of Generation Strategies on Text Quality: In the table, FS stands for Few-Shot, and Pipeline is based on the Automated Dataset Generation Pipeline we proposed in Section 4 and our proposed dataset in Section 3. The prompt used in this experiment can be found in Appendix A.6. The purpose is to demonstrate that, using the same prompt, our pipeline achieves the highest text generation quality that meets our task requirements.
5.3.1 Six-Dimensional Analysis of Depression task
Research has demonstrated that structured knowledge enhances the benefits of integrating structured data into LLMs Moiseev et al. (2022), thereby improving output quality. CNSD Gold and CNSD Silver are datasets with fine-grained structured annotations across six dimensions to summarize and analyze social media users. Fine-tuning on these datasets should reduces textual hallucinations and achieves text generation quality comparable to the other much larger model.
Table 3 shows the comparisons between different generation strategies in terms of the quality of the produced text. Notably, our Pipeline achieves the highest scores across all three metrics:BERTScore (0.791), ROUGE-1 (0.478), and BLEU (0.288), indicating superior semantic alignment, lexical overlap, and n-gram precision, respectively. In contrast, with the FS (few-shot) strategies, all models: Qwen2.5 14B, GPT-4o, and DeepSeek R1 671B, yield comparatively lower performance on each metric. Even the large-scale DeepSeek R1 671B model does not surpass the Pipeline approach, suggesting that the proposed Pipeline method effectively leverages its generation procedure to produce higher-quality text that more closely matches the reference, outperforming both smaller and larger models under the few-shot setting.
Additionally, we also report the comparison of generation quality between the original model, Qwen2.5 14B Gold, and Qwen2.5 14B Silver. (Table 5)
For automatic metrics, the original model achieved a Bert-Score of 0.649, which increased to 0.764 for Gold (+16.9%) and 0.787 for Silver (+21.5%), showing stronger semantic alignment. Rouge-1 rose from 0.076 to 0.217 for Gold (+172%) and 0.218 for Silver (+172%), indicating improved lexical overlap. Bleu increased from 0.075 to 0.186 for Gold (+149%) and 0.237 for Silver (+218%), reflecting better n-gram precision and structural quality.
The human evaluation was conducted by two linguistic experts across accuracy, coverage, and hallucination. Compared to the original Qwen2.5 14B, both Silver and Gold showed clear improvements. Silver improved by about 33.4% in accuracy, 37.1% in coverage, and 15.5% in hallucination. Gold achieved even stronger gains, with 42.9% in accuracy, 58.0% in coverage, and 26.8% in hallucination. Overall, Gold performed best in human evaluation, while Silver, though slightly behind, still showed significant gains over the original and reached levels close to Gold.
In summary, Gold achieved the strongest results in human evaluation, while Silver performed better in automatic metrics. Both models clearly outperformed the original, showing that fine-tuning with our dataset or pipeline can significantly enhance generation quality at semantic, lexical, and structural levels.
Model/Test Dataset Swdd Origin Wu3d Ours (Gold)
Acc F1 Acc F1 Acc F1
RoBERTa Chinese (trained by swdd origin) 0.89 0.90 0.76 0.80 0.88 0.87
RoBERTa Chinese (trained by wu3d) 0.88 0.87 0.90 0.91 0.82 0.78
Bert based Chinese (trained by swdd origin) 0.89 0.90 0.76 0.80 0.85 0.82
Bert based Chinese (trained by wu3d) 0.88 0.87 0.90 0.91 0.83 0.80
StructBERT-mental (trained by swdd origin) 0.81 0.84 0.69 0.76 0.85 0.86
StructBERT-mental (trained by wu3d) 0.85 0.85 0.92 0.92 0.83 0.80
Llama3-8b 0.86 0.86 0.79 0.79 0.60 0.63
Glm4-9b-chat 0.76 0.81 0.77 0.81 0.79 0.82
Qwen2.5 7b 0.88 0.89 0.84 0.85 0.94 0.94
GPT-4o-Mini 0.65 0.74 0.65 0.74 0.53 0.68
GPT-4o 0.92 0.92 0.91 0.92 0.83 0.86
DeepSeek-R1 671b 0.86 0.87 0.89 0.90 0.81 0.84
Table 4: Model Classification Performance Comparison on Different Test Datasets.
Model Bert-Score Rouge1 Bleu H.acc H.cov H.hallu
Qwen2.5 14B 0.649 0.076 0.075 0.583 0.574 0.657
Qwen2.5 14B Silver 0.787 0.218 0.237 0.778 0.787 0.759
Qwen2.5 14B Gold 0.764 0.217 0.186 0.833 0.907 0.833
Table 5: Comparison of text quality across different models. ’Qwen2.5 14B Gold’ denotes the model fine-tuned on our proposed dataset, while ’Qwen2.5 14B Silver’ denotes the model fine-tuned using data generated from the pipeline strategy described in Table 3, with further details in Section 4. The table reports three human evaluation metrics: acc (accuracy of the generated text), cov (content coverage), and hallu (hallucination). Higher values across all metrics indicate better generation performance.
5.3.2 Classification task
In the depression classification task, Qwen2.5 14B Silver achieves the best performance compared to other models, with an accuracy of 0.944 and an F1 score of 0.941, even surpassing Qwen2.5 14B Gold, which was fine-tuned on human-annotated data. Furthermore, it outperforms large-scale models such as GPT-4o (accuracy = 0.917, F1 score = 0.923) and DeepSeek R1 671B (accuracy = 0.861, F1 score = 0.872). This result provides evidence for the effectiveness of our generation pipeline and the high quality of the data produced through this pipeline. The details of the experimental results are in the Table 6.
Model Accuracy Recall Precision F1
Qwen2.5 14B 0.889 0.944 0.850 0.894
Qwen2.5 14B_FT_Silver 0.944 0.889 1.000 0.941
Qwen2.5 14B_FT_Gold 0.861 0.889 0.842 0.864
GPT-4o 0.917 1.000 0.857 0.923
GPT-4o Mini 0.667 1.000 0.600 0.750
DeepSeek R1 671B 0.861 0.944 0.801 0.872
DeepSeek R1 Distill-14B 0.889 1.000 0.818 0.900
Table 6: Comparison of Models for Generated Dataset Quality: Application to Classification Tasks.
5.4 Task II: Structured Summarization and Analysis Generation
The proposed dataset effectively supports research on generative models in the field of internet depression risk analyses, demonstrating particular value for user depressive mental state summarization tasks. Due to the absence of comparable open-source datasets, we conducted preliminary evaluations on mainstream generative models using our annotated dataset containing 233 user-level samples. The detailed results are demonstrated in the table 7. We used the basic prompt as follows:
Please carefully read the following multiple texts and then answer the question: Based on these texts, does this user exhibit a depressive mood? If the user exhibits depressive mood, answer ’Yes’; if the user does not exhibit depressive mood, answer ’No’. Then, provide your step-by-step analysis process.
Model Bert-Score Rouge1 Bleu
Llama3-8b $0.654$ $0.194$ $0.122$
Glm4-9b-chat $0.694$ $0.208$ $0.140$
Qwen2.5-7b $0.680$ $0.208$ $0.093$
DeepSeek-R1-671B $0.698$ 0.454 $0.125$
GPT-4o-mini $0.696$ $0.166$ $0.126$
GPT-4o 0.710 $0.346$ 0.152
Table 7: Performance Comparison on Depression Risk Analysis Summarization Generation Task.
The results reveal variations across models in the task of risk analysis summarization generation:
Bert-Score: Narrow range (0.65-0.71) with GPT-4o leading at 0.71, indicating superior semantic understanding. Rouge1: DeepSeek-R1-671B achieves remarkable 0.45 score, suggesting 10% improvement over second-best performer. Bleu: Maximum 15.17 from GPT-4o vs minimum 9.36 from Qwen2.5-7b, reflecting lexical alignment differences.
5.5 Task III: Classification Experimentations
In the depression risk binary classification task, we used three datasets, including the original SWDD dataset Cai et al. (2023) and the WU3D dataset Wang et al. (2020). Both of these datasets were randomly sampled with 200 negative users and 200 positive users for the experiment. Details provided in the table 4.
In the classification experiments, in addition to the generative models mentioned in the section A.1, we also used models based on the BERT architecture with the details on A.7.
6 Conclusion
We introduced CNSocialDepress, a new dataset for depression risk analysis using Chinese social media texts. Uniquely, it is the first dataset to include both user-level and text-level labels, combining binary depression indicators with fine-grained six-dimensional analysis annotated by mental health professionals. We also developed an automated data generation pipeline to ease manual annotation and support language model fine-tuning for classification and summarization tasks. We believe CNSocialDepress will prove valuable across a range of applications, advancing early detection and intervention strategies in mental health.
7 Limitations
While CNSocialDepress provides a valuable resource for depression risk analyses in Chinese social media, several limitations should be noted. First, the dataset is sourced exclusively from Weibo, potentially introducing platform-specific demographic biases (e.g., underrepresentation of rural or elderly populations).
Second, linguistic diversity, including dialectal variations and metaphorical expressions common in Chinese depressive discourse, remains insufficiently captured and may limit cross-regional generalizability. Third, although manual expert annotation ensures high quality, it constrains scalability, resulting in a relatively limited dataset size.
Additionally, the current structured analysis focus primarily on depressive symptoms, omitting comorbid mental health conditions (e.g., anxiety) and contextual social factors. Future work will expand data collection to multi-platform sources, integrate dynamic lexicon updates for emerging expressions, and adopt hybrid annotation frameworks (e.g., expert-guided crowdsourcing) to enhance coverage and efficiency. Privacy-preserving techniques and longitudinal data tracking will further strengthen ethical and practical utility.
8 Ethic Statement
This dataset is built upon the SWDD Cai et al. (2023) benchmark released in early 2023. The original data comes from user-level text on the Weibo platform and was annotated by a team of psychology experts according to the DSM-5 symptom standards.
During our work on the re-annotation process, we applied a second round of deep anonymization to meet privacy requirements. This involved systematically removing personal information, geographic locations, and potential identifiers from the original text, ensuring compliance with data protection protocols Benton et al. (2017) and GDPR (General Data Protection Regulation [^8] EUR-Lex GDPR Regulation). We also invited experienced researchers specializing in depression-related psychology to provide professional annotations based on their expertise.
This dataset is mainly used for building early-warning models for depression risk. Through multi-dimensional feature analysis, it detects potential risk signals and can be used to automatically generate structured psychological assessment reports, offering technical support for digital mental health services. It is important to note that this study is strictly limited to the realm of mental health assistance, and any results should only serve as references for professional diagnosis or as a digital tool for self-assessment.
At present, none of the algorithmic models can replace in-person psychiatric evaluations or deliver pathological diagnoses. If a system built using this dataset flags high-risk cases, or if individuals already show persistent mood disturbances or impaired social functioning, they should seek professional medical attention promptly. Mental health issues require clinical expertise, and any online self-assessment tool cannot substitute for in-person medical diagnosis.
Appendix A Appendix
A.1 Baseline Models
  • DeepSeek-R1 is a large-scale model specialized in logical reasoning,This model excels in complex reasoning tasks such as mathematics and programming while maintaining strong overall performance on general tasks. Chosen for its specialized excellence in complex logical reasoning while maintaining strong general task performance, potentially beneficial for in-depth structured psychological analysis and nuanced depression risk detection.
  • DeepSeek R1 Distill-14B is a lightweight reasoning model derived from Qwen2.5-14B through Knowledge Distillation. Chosen for its lightweight reasoning capabilities, aiming to retain the advantages of larger models in complex tasks while reducing computational cost, making it suitable for efficient depression risk analysis.
  • Qwen2.5-14B is part of Alibaba’s Qwen series of large language models. Chosen for its strong performance in long-form text generation and instruction-following, relevant for generating structured psychological analyses from Chinese social media posts.
  • GPT-4o is a multimodal large language model released by OpenAI. Chosen for its state-of-the-art multimodal and multilingual processing capabilities, including Chinese, and its improved overall performance and response speed, making it a powerful baseline for depression detection and psychological analysis.
  • GPT-4o-Mini is a lightweight AI model developed by OpenAI to provide an efficient and cost-effective AI solution. Chosen as a resource- and cost-efficient model with robust text understanding and generation, suitable for less computationally demanding depression risk identification and basic psychological analysis tasks.
  • Llama3-8B-Chinese-Chat, derived from Meta-Llama-3-8B, is an instruction-tuned model optimized for Chinese dialogue tasks. It has been fine-tuned to process both Chinese and English inputs. Chosen for its optimization for Chinese dialogue tasks and ability to handle extended conversations, making it well-suited for analyzing conversational social media posts for depression risk and generating coherent psychological insights in Chinese.
A.2 20 Randomly Generated Expression for the Positive Users
  1. 1.
    Carefully read the complete user Weibo text below and extract key information across six specified dimensions according to the content. Also pay attention to the user’s depressive state, ensuring that each dimension’s explanation is consistent with this state.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - For each dimension, if evidence exists in the text, provide the text index (e.g., text_xx, 文本_xx) and a brief note (e.g., text_xx[sad]). - If no evidence is found for a dimension, state clearly that no relevant evidence was found.
    Note: Only output text index and explanation, not the actual text.
  2. 2.
    Read the entire user Weibo text below and, based on its content, extract core information under the following six dimensions. Also consider the user’s depressive status, ensuring that each dimension’s description matches this state.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - For each dimension, if relevant information is detected, provide the text index (e.g., text_xx, 文本_xx) and a short note (e.g., text_xx[unhappy]). - For dimensions without evidence, explicitly state no relevant evidence found.
    Note: Output should include only text index and explanation, not the actual text.
  3. 3.
    Read through the complete user Weibo text below and extract key information across six dimensions based on its content. Pay close attention to the user’s depressive state, ensuring that the explanations are consistent with it.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - If relevant evidence is found in a dimension, provide the text index (e.g., text_xx, 文本_xx) and a short note (e.g., text_xx[sorrow]). - If no evidence is found, indicate explicitly.
    Note: Provide only indices and explanations, not the original content.
  4. 4.
    Carefully read the full Weibo text below and extract key evidence across six dimensions according to the text. At the same time, take into account the user’s depressive state, ensuring that explanations for each dimension align with this state.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - For each dimension, if there is corresponding evidence, output the text index (e.g., text_xx, 文本_xx) and a related note (e.g., text_xx[frustrated]). - If no evidence is found, state that no evidence exists.
    Note: Only provide text indices and notes, not the raw text.
  5. 5.
    Please read the full Weibo text below and extract key information from six dimensions. At the same time, pay attention to the user’s depressive state, ensuring that your explanations match this state.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - For each dimension, if evidence is detected, output the text index (e.g., text_xx, 文本_xx) and a short description (e.g., text_xx[sadness]). - If no evidence is found, specify that no evidence was found.
    Note: Only output text indices and explanations, not the original text.
  6. 6.
    Read the complete Weibo text below carefully and extract key information across six dimensions based on the text. Pay attention to the user’s depressive state and ensure that your descriptions are consistent with it.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - If relevant evidence is found, provide the text index (e.g., text_xx, 文本_xx) and a brief description (e.g., text_xx[low mood]). - If no evidence is found, clearly state that no evidence exists.
    Note: Only provide text indices and explanations, not the original text.
  7. 7.
    Please read through the complete user Weibo text below and extract key information across six dimensions based on its content. Pay attention to the user’s depressive state, ensuring that each explanation is aligned with this state.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - For each dimension, if evidence exists, provide the text index (e.g., text_xx, 文本_xx) and a short note (e.g., text_xx[loss]). - If no evidence is found, clearly state no relevant evidence.
    Note: Only provide text indices and notes, not the original text.
  8. 8.
    Read the Weibo text below and extract key evidence across six dimensions according to the content, while considering the user’s depressive state to ensure explanations are consistent.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - If a dimension contains relevant evidence, provide the text index (e.g., text_xx, 文本_xx) and a short note (e.g., text_xx[low mood]). - If no evidence is found, state explicitly that no evidence exists.
    Note: Output only text indices and explanations, not the original text.
  9. 9.
    Read the complete Weibo text carefully and extract key evidence across six dimensions based on the text. Pay attention to the user’s depressive state, ensuring that explanations match this state.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - For each dimension, if relevant evidence is found, output the text index (e.g., text_xx, 文本_xx) and a brief explanation (e.g., text_xx[depressed]). - If no evidence is found, state explicitly that there is none.
    Note: Provide only text indices and notes, not the original text.
  10. 10.
    Read through the entire Weibo text and extract core information across six dimensions according to the content. Pay attention to the user’s depressive state and ensure that the descriptions are consistent with this state.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - For each dimension, if relevant evidence exists, output the text index (e.g., text_xx, 文本_xx) and a short explanation (e.g., text_xx[pessimistic]). - If no evidence is found, state clearly that no evidence exists.
    Note: Only output text indices and explanations, not the original content.
  11. 11.
    Please read the user Weibo text below completely and extract essential information across six dimensions. Ensure that the explanations for each dimension align with the depressive state given.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - If relevant evidence is detected, list the text index (e.g., text_xx, 文本_xx) and a short description (e.g., text_xx[melancholy]). - If no evidence is found, specify that no evidence was detected.
    Note: Output should only contain indices and notes, not the actual text.
  12. 12.
    Read carefully the full Weibo text provided and extract key evidence from six dimensions. Pay attention to the user’s depressive state, ensuring that outputs remain consistent with this state.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - If a dimension contains evidence, output the text index (e.g., text_xx, 文本_xx) and a concise note (e.g., text_xx[disheartened]). - If no evidence exists, specify so.
    Note: Output indices and explanations only, not original content.
  13. 13.
    Please go through the complete Weibo text below and identify critical information across six dimensions. Ensure the notes correspond with the user’s depressive condition.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - Where evidence is found, list the text index (e.g., text_xx, 文本_xx) with a short explanation (e.g., text_xx[downhearted]). - If no evidence is detected, indicate explicitly.
    Note: Provide only text indices and notes, excluding original text.
  14. 14.
    Read carefully the following Weibo text and extract evidence across six preset dimensions. Make sure the explanation for each aligns with the depressive state.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - If evidence exists for a dimension, output the text index (e.g., text_xx, 文本_xx) and a short remark (e.g., text_xx[low mood]). - If no evidence is found, specify clearly.
    Note: Output only indices and remarks, not the raw text.
  15. 15.
    Read through the following Weibo text and extract major evidence across six given dimensions. Ensure the output is consistent with the depressive state specified.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - If evidence is found, provide the text index (e.g., text_xx, 文本_xx) with a brief note (e.g., text_xx[hopeless]). - If no evidence appears, state clearly.
    Note: Only provide indices and notes, not original text.
  16. 16.
    Please carefully review the Weibo text below and extract key information from six preset dimensions. All notes must align with the depressive condition of the user.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - If there is evidence, output the corresponding text index (e.g., text_xx, 文本_xx) and a short remark (e.g., text_xx[emotional slump]). - If not found, specify clearly that none exists.
    Note: Only include indices and notes, no original content.
  17. 17.
    Please go through the complete Weibo text and extract the critical information across six given dimensions. Make sure your notes are consistent with the depressive state provided.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - For each dimension, if evidence is found, provide the text index (e.g., text_xx, 文本_xx) with a concise explanation (e.g., text_xx[feeling low]). - If no evidence exists, indicate so.
    Note: Only output indices and explanations, not actual text.
  18. 18.
    Carefully read the Weibo text provided and extract important evidence across six specified dimensions. Keep the explanations aligned with the depressive state.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - If a dimension contains evidence, state the text index (e.g., text_xx, 文本_xx) and a short remark (e.g., text_xx[upset]). - If not, state that no evidence is found.
    Note: Output only indices and remarks, not text.
  19. 19.
    Read the Weibo text completely and extract relevant evidence from six dimensions. Ensure each dimension’s note is consistent with the depressive state given.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - If relevant evidence is detected, provide the text index (e.g., text_xx, 文本_xx) and short description (e.g., text_xx[distressed]). - If not found, state clearly no evidence.
    Note: Only indices and notes, exclude text itself.
  20. 20.
    Please read through the following full Weibo text and extract essential evidence across six preset dimensions. Ensure all notes are consistent with the depressive state.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - If evidence is found, provide the text index (e.g., text_xx, 文本_xx) and a short explanation (e.g., text_xx[worry]). - If no evidence is detected, state that explicitly.
    Note: Only indices and explanations, not original text.
A.3 20 Distinct Label Expressions for the Negative Users
  1. 1.
    "This text does not contain any expressions of negative emotions, nor does it involve any dimensions of depression."
  2. 2.
    "There are no descriptions related to negative emotions in the text, so it does not fit any depression-related dimensions."
  3. 3.
    "This text does not include expressions of negative emotions, and therefore it is not classified under any depression dimensions."
  4. 4.
    "The content does not reflect negative emotions and does not belong to any depression indicators."
  5. 5.
    "This passage contains no expressions of negative emotions and is not applicable to any depression dimensions."
  6. 6.
    "There are no descriptions of negative emotions in the text, so it is not classified under any depression-related dimensions."
  7. 7.
    "This content does not exhibit negative emotions and therefore does not involve any depression dimensions."
  8. 8.
    "The text does not contain any expressions related to negative emotions, nor does it meet the criteria for depression dimensions."
  9. 9.
    "There are no signs of negative emotions in this text, so it does not belong to any depression dimension."
  10. 10.
    "The text does not describe negative emotions and does not cover any depression-related dimensions."
  11. 11.
    "This passage does not contain any expressions related to negative emotions and is not within the scope of depression dimensions."
  12. 12.
    "The content does not include expressions of negative emotions, so it does not fall into any depression dimensions."
  13. 13.
    "There are no expressions of negative emotions in the text, nor does it meet the criteria for depression dimensions."
  14. 14.
    "This text does not exhibit any negative emotions and does not involve any depression dimensions."
  15. 15.
    "This content does not include descriptions of negative emotions and therefore is not classified under any dimensions of depression."
  16. 16.
    "The text does not show any negative emotions and is not applicable to depression-related dimensions."
  17. 17.
    "No expressions of negative emotions are seen in this passage, so it does not meet any depression dimension standards."
  18. 18.
    "There are no expressions of negative emotions in the text, so it does not belong to any depression dimension."
  19. 19.
    "This text does not display any expressions related to negative emotions and does not cover the dimensions of depression."
  20. 20.
    "The content lacks descriptions of negative emotions and is therefore not classified under any depression dimensions."
A.4 Instructions for the Fine-Tuning Process
  • Complete Instruction = Task Instruction + Supplementary Instruction
Task Instruction:
This task involves the following 6 dimensions of depression:
  • Potential External Causes of Depression (Secondary Judgment Criterion)
  • Depression-Related Clinical Symptoms (Primary Judgment Criterion)
  • Depression-Related Language Expression Patterns (Secondary Judgment Criterion)
  • Depression-Related Medical Expressions (Primary Judgment Criterion)
  • Depressive Psychological State (Primary Judgment Criterion)
  • Negative Emotions (Secondary Judgment Criterion)
Supplementary Instructions (randomly select one):
  1. 1.
    "Please read the following text segment and determine whether it contains any of the above 6 expressions of depressive emotion, and provide a brief explanation; if none is present, please indicate that the text does not belong to any depression dimension."
  2. 2.
    "Read the following text and analyze whether it demonstrates any one of the aforementioned 6 depressive emotion expression dimensions, and provide a brief explanation; if such expression is absent, please indicate that the text does not correspond to any dimension."
  3. 3.
    "Please carefully read the following text segment and confirm whether any one of the 6 depressive emotion expression dimensions mentioned above exists, and include a brief explanation; if not, please state that the text does not belong to any dimension."
  4. 4.
    "Please read the following content and determine whether it embodies any one of the above 6 depressive emotion expression dimensions, and provide a brief explanation; if not, please indicate that the text does not cover any depression dimension."
  5. 5.
    "Read the following text segment and determine whether it contains any one of the 6 depressive emotion expressions mentioned above, and provide a concise explanation; if not, please state that the text does not belong to any depression dimension."
  6. 6.
    "Please read the following text and determine whether any one of the above 6 depressive emotion expression dimensions appears, and include a brief explanation; if not, please indicate that the text does not meet any dimension."
  7. 7.
    "Read the following text and check whether any one of the above 6 depressive emotion expression dimensions is presented, and provide a brief explanation; if not, please state that the text does not belong to any dimension."
  8. 8.
    "Please review the following text segment and determine whether any one of the above 6 depressive emotion expression dimensions exists, and provide a brief explanation; if not, then indicate that the text does not belong to any dimension."
  9. 9.
    "Read the following text and confirm whether any one of the above 6 depressive emotion expression dimensions is embodied, and provide a brief explanation; if not, please state that the text does not involve any depression dimension."
  10. 10.
    "Please read the following content and determine whether it contains any one of the 6 depressive emotion expression dimensions mentioned earlier, and provide a brief explanation; if not, please indicate that the text does not belong to any dimension."
  11. 11.
    "Read the following text segment and verify whether it demonstrates any one of the above 6 depressive emotion expression dimensions, and include a brief explanation; if not, please state that the text does not meet any dimension."
  12. 12.
    "Please read the following text and check whether it exhibits any one of the above 6 depressive emotion expression dimensions, and provide a brief explanation; if not, please indicate that the text does not cover any depression dimension."
  13. 13.
    "Read the following text and confirm whether any one of the 6 depressive emotion expressions mentioned above exists, and include a brief explanation; if such a feature is absent, please state that the text does not belong to any dimension."
  14. 14.
    "Please carefully read the following text segment and determine whether it embodies any one of the above 6 depressive emotion expression dimensions, and provide a concise explanation; if not, please indicate that the text does not correspond to any dimension."
  15. 15.
    "Read the following text segment and confirm whether it contains any one of the above 6 depressive emotion expression dimensions, and provide a brief explanation; if not, please state that the text does not involve any depression dimension."
  16. 16.
    "Please read the following text segment and determine whether any one of the 6 depressive emotion expression dimensions mentioned above appears in the text, and include a brief explanation; if not, please indicate that the text does not belong to any dimension."
  17. 17.
    "Read the following content and check whether it contains any one of the above 6 depressive emotion expression dimensions, and provide a brief explanation; if not, please state that the text does not belong to any depression dimension."
  18. 18.
    "Please read the following content and determine whether any one of the above 6 depressive emotion expression dimensions exists, and provide a brief explanation; if not, please indicate that the text does not involve any dimension."
  19. 19.
    "Read the following text segment and confirm whether it demonstrates any one of the above 6 depressive emotion expression dimensions mentioned above, and provide a brief explanation; if not, please state that the text does not meet any dimension."
  20. 20.
    "Please read the following text and determine whether it contains any one of the above 6 depressive emotion expression dimensions, and provide a concise explanation; if not, please indicate that the text does not belong to any depression dimension."
A.5 Module II Prompt
Figure 3: Module II Prompt.
A.6 Prompt for Table 3 On Section 5.3.1
Prompt is as follows: You are a professional in psychology and text sentiment analysis. Please determine whether the user is at risk of depression based on the following six dimensions:
  • Negative emotions (secondary criterion)
  • Depressive psychological state (primary criterion)
  • Depression-related clinical symptoms (primary criterion)
  • Potential external factors causing depression (secondary criterion)
  • Depression-related medical expressions (primary criterion)
  • Depression-related language expression patterns (secondary criterion)
For each dimension, please cite the corresponding text reference and provide a concise explanation.
[Output Data Example]
  • Answer: “Yes” or “No.”
  • Briefly explain the manifestations in each of the six dimensions, and cite the corresponding text reference.
  • In conclusion, “The user is at risk of depression” or “The user is not at risk of depression.”
A.7 Bert type models
BERT-base-chinese [^9] https://huggingface.co/google-BERT/BERT-base-chinese is built on the BERT-Base architecture Devlin et al. (2019) and is pre-trained on extensive Chinese corpora. It adopts masked language modeling (MLM) and next sentence prediction (NSP) to learn deep bidirectional representations of Chinese text at the character level.
Chinese-RoBERTa-wwm-ext Cui et al. (2019) leverages RoBERTa Liu et al. (2019) with a whole word masking (WWM) approach. Instead of masking individual characters, it masks entire words during MLM, leading to a richer understanding of context. The extended pre-training (ext) on large Chinese corpora further refines its effectiveness in various NLP tasks.
StructBERT-mental Wang et al. (2019) is based on StructBERT, which incorporates linguistic structures into its pre-training. By focusing on word and sentence structure, and adding specific pre-training for mental health texts, it improves performance in identifying and categorizing mental health-related content.